the length of a swimming pool is 12 feet longer than 1/2 its width. the perimeter of the pool is 60 feet. find the length of the pool?
Answer by checkley77(12844)

L=W/2+12
2L+2W=60
2(W/2+12)+2W=60
W+24+2W=60
3W=60-24
3W=36
W=36/3
W=12 FT. FOR THE WIDTH.
L=12/2+12
L=6+12=18 FT. FOR THE LENGTH.
PROOF:
2*18+2*12=60
36+24=60
60=60